Session: Oct 2010 (session 3)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Susana Garcia-Espada at Thu Oct 21 15:05:33 2010) Minor failures: RFI
Strong interference at BBC07 lower band (only RCP?) or agc is not working properly for BBC7 lower band?

Medicina (Andrea Orlati at Thu Oct 21 16:08:08 2010) Minor failures:
Vc02 unlock alarm

Effelsberg (Uwe Bach at Fri Oct 22 10:31:23 2010) Success:
No known problems!

Torun (K.Borkowski at Fri Oct 22 11:00:52 2010) Success:
Off source recording between 14:06 and 14:41 UTC due to an outage (therefore also the third ftp scan missed).

Westerbork (Geert Kuper at Fri Oct 22 14:21:18 2010) Success:

Onsala (Michael Lindqvist at Sun Oct 24 06:21:39 2010) Success:
No known problems.